If G be the the geometric mean of x and y then 1/G^2-x^2 +1/G^2-y^2?

 

Grade:12

1 Answers

Arun
6 years ago
G^2 = XY 
SO 1/(G²-x²)+1/(G²-y²) 
= 1/(XY-X^2) + 1/(XY-Y^2) 
= 1/X* 1/(Y-X)+1/Y* 1/(X-Y) 
= 1/(x-y)[1/y - 1/x] 
=1/(x-y) *[(x-y)/xy] 
= 1 / xy 
= 1 / G^2